# Cost Efficient Fairness Audit Under Partial Feedback

### v1

@@ 4.1 RS-Audit: Rejection Sampling based Audit
$\mathbb{P}[f=1\mid Y=y,A=a]=\frac{\mathbb{P}[f=1,Y=y\mid A=a]}{\mathbb{P}[Y=y\mid A=a]}=\frac{p_{y\mid a}}{q_{y\mid a}}~.$
@@
1: Input $\varepsilon,\delta\in(0,1)$ and past database $D$, then set $\tau={576\log(8\lvert\mathcal{A}\rvert/\delta)}/{\varepsilon^{2}}$ and $\widehat{\Delta}=0$.
@@
4:   $N\leftarrow\texttt{OnlineSample}(\tau,y,a)$, $N^{\prime}\leftarrow\texttt{PastSample}(D,\tau,y,a)$.
5:   Set $\widehat{q}_{y\mid a}=\frac{\tau}{N}$ and $\widehat{p}_{y\mid a}=\frac{\tau}{N^{\prime}}$. Update $\widehat{\Delta}\leftarrow\max\left\{\widehat{\Delta},\left\lvert\frac{\widehat{p}_{y\mid a}}{\widehat{q}_{y\mid a}}-\frac{\widehat{p}_{y\mid a^{\prime}}}{\widehat{q}_{y\mid a^{\prime}}}\right\rvert\right\}$
@@
Although RS-Audit and baseline (Algorithm 3) are superficially similar, these are two conceptually different algorithms. The novel insight of not observing the true labels for individuals with $A\neq a$ when trying to estimate $q_{y\mid a}$ exploits the structure of the audit cost under partial feedback where cost accrues only when the auditor requests the true label. Moreover, we crucially identify the right amount of information needed to estimate $\mathbb{P}[f=1\mid Y=y,A=a]$, namely, samples of true labels from the distribution $\mathbb{P}[Y\mid A=a]$. Combining these two observations lead to RS-Audit, which is nearly optimal in sample complexity (see Theorem 3). Moreover, due to these observations, the gap in sample complexity of baseline over RS-Audit is roughly $\widetilde{O}\left(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{\mathbb{P}[f=0,A\neq a]}{q_{y,a}\varepsilon^{2}}\right)$ and the difference in audit cost is roughly $\widetilde{O}\left(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{c_{feat}\mathbb{P}[f=0,A\neq a]+c_{lab}\mathbb{P}[Y=0,f=0,A\neq a]}{q_{y,a}\varepsilon^{2}}\right)$.
@@ 5 Fairness Audit under Mixture Model
Recall that in the mixture model, given $A=a$, the true label is sampled according to $Y\sim Ber(q_{1\mid a})$ and given both $A=a$ and $Y=y$, the feature of the individual is sampled according to $X\sim{\mathcal{E}}_{\theta^{*}_{y,a}}$ (see Section 2 and 3). Also, we denote $\mathbf{c}\coloneqq(L,\beta,\kappa,\alpha)$, which are some known quantities of the problem. Now we describe Exp-Audit (Algorithm 2) that solves fairness audit under partial feedback in this model with a significantly lower audit cost than the blackbox case.
Algorithm 2 Exp-Audit: Fairness Audit Algorithm for Mixture Model
@@
1: Set $\varepsilon^{\prime}={1}/{10}$, $\tau^{\prime}={4\log(14\lvert{\mathcal{A}}\rvert/\delta)}/{\varepsilon^{\prime 2}}$, $\tau={576\log(14\lvert{\mathcal{A}}\rvert/\delta)}/{\varepsilon^{2}}$ and $\widehat{\Delta}=0$.
@@
5:   $N^{\prime}\leftarrow\texttt{OnlineSample}(\tau^{\prime},y,a)$, $N\leftarrow\texttt{PastSample}(D,\tau,y,a)$
@@
7:  Construct MAP oracle: $M_{a}(x)=\text{argmax}_{y\in\{0,1\}}\widetilde{q}_{y\mid a}\mathcal{E}_{\widehat{\theta}_{y,a}}(x).$
8:  Let $\widetilde{q}_{m,a}=\widetilde{q}_{1\mid a}\wedge\widetilde{q}_{0\mid a},~R_{0}\!=\!R_{1}\!=\!0$. Observe $R=\frac{3430\log(12\lvert{\mathcal{A}}\rvert/\delta)}{\widetilde{q}_{m,a}\varepsilon^{2}}$ features $X$ with A=aA\!=\!a and set $R_{M_{a}(X)}\leftarrow R_{M_{a}(X)}+1$ for each $X$.
9:  Set $\widehat{q}_{y\mid a}={R_{y}}/{R}$ and update $\widehat{\Delta}\leftarrow\max\left\{\widehat{\Delta},\left\lvert\frac{\widehat{p}_{y\mid a}}{\widehat{q}_{y\mid a}}-\frac{\widehat{p}_{y\mid a^{\prime}}}{\widehat{q}_{y\mid a^{\prime}}}\right\rvert\right\}$
@@
Exp-Audit first splits the past database of positively classified individuals into subsets $D_{y,a}$, $a\in{\mathcal{A}}$, $y\in\{0,1\}$, so that $D_{y,a}$ only contains individuals with $A=a$ and $Y=y$. Note that samples in $D_{y,a}$ are iid samples of $X\sim{\mathcal{E}}_{\theta^{*}_{y,a}}$ truncated by $f$, i.e., $D_{y,a}$ contains only those samples from ${\mathcal{E}}_{\theta^{*}_{y,a}}$ for whom $f=1$. On every subset $D_{y,a}$, Exp-Audit applies the subroutine $\texttt{TruncEst}(D_{y,a},f,\frac{\varepsilon}{2\beta},\frac{\delta}{14\lvert{\mathcal{A}}\rvert},\mathbf{c},\Theta)$ (an adaptation of LWZ (23), described in Appendix B.2). From $D_{y,a}$ this subroutine learns an estimate $\widehat{\theta}_{y,a}$ of $\theta^{*}_{y,a}$ such that $\lVert\widehat{\theta}_{y,a}-\theta^{*}_{y,a}\rVert\leq\frac{\varepsilon}{2\beta}$ with probability at least $1-\frac{\delta}{14\lvert{\mathcal{A}}\rvert}$ if $\lvert D_{y,a}\rvert\geq\widetilde{\Omega}(\frac{d}{\varepsilon^{2}})$ (which can be assumed true without loss of generality [^2] The auditor can wait without any extra audit cost till this condition is met which happens with high probability in finite time due to the positivity assumption $\mathbb{P}[f=1\mid Y=y,A=a]>0$.). Next, the algorithm uses the rejection-sampling based approach similar to RS-Audit (line 5) to obtain $\widehat{p}_{y\mid a}\in(1\pm\frac{\varepsilon}{12})p_{y\mid a}$ and $\widetilde{q}_{y\mid a}\in(1\pm\varepsilon^{\prime})q_{y\mid a}$, where $\varepsilon^{\prime}=\frac{1}{10}$ is a constant. Note that the cost incurred in estimating $\widetilde{q}_{y,a}$ is therefore $\widetilde{O}(1)$. This is a crucial difference of Exp-Audit with RS-Audit which estimates $q_{y\mid a}$ upto a multiplicative error of $1\pm\varepsilon$ and incurs $\widetilde{O}(\frac{1}{\varepsilon^{2}})$ cost.
@@
After this, for every $a\in{\mathcal{A}}$, Exp-Audit constructs a maximum a posteriori (MAP) oracle for the mixture model $M_{a}:{\mathcal{X}}\to\{0,1\}$, $M_{a}(X)=\operatorname*{arg\,max}_{y\in\{0,1\}}\widetilde{q}_{y\mid a}{\mathcal{E}}_{\widehat{\theta}_{y,a}}(X)$. In other words, for $A=a$ and feature $X$, the MAP oracle returns which is the most likely true label $Y$. Next, for every $a\in{\mathcal{A}}$, with $\widetilde{q}_{m,a}=\widetilde{q}_{0\mid a}\wedge\widetilde{q}_{1\mid a}$, the algorithm observes only the features $X$ for $R=\widetilde{O}(\frac{1}{\widetilde{q}_{m,a}\varepsilon^{2}})$ individuals from group $A=a$ and uses the MAP oracle $M_{a}$ to assign a proxy label to each $X$. With $R_{y}$ as the number of individuals with proxy label $y$, Exp-Audit computes $\widehat{q}_{y\mid a}=\frac{R_{y}}{R}$. Under the separation condition (1), we have $\widehat{q}_{y\mid a}\in(1\pm\frac{\varepsilon}{12})q_{y\mid a}$ (see Appendix B.1). Using $\widehat{p}_{y\mid a}$ and $\widehat{q}_{y\mid a}$, Exp-Audit sets $\widehat{\Delta}=\max_{y\in\{0,1\},a,a^{\prime}\in{\mathcal{A}}}\left\lvert\frac{\widehat{p}_{y\mid a}}{\widehat{q}_{y\mid a}}-\frac{\widehat{p}_{y\mid a^{\prime}}}{\widehat{q}_{y\mid a^{\prime}}}\right\rvert$ and returns $\mathsf{UNFAIR}$ if $\widehat{\Delta}>\frac{\varepsilon}{2}$, else $\mathsf{FAIR}$.
@@
The cost incurred by Exp-Audit is significantly lower than that of RS-Audit (see Theorems 2 and 4). In the guarantee of Exp-Audit (Theorem 4), the term involved with $c_{lab}$, the cost of requesting a negative label ($Y=0$), is a problem-dependent constant and independent of $\varepsilon$. In most applications, $c_{lab}\gg c_{feat}$ and therefore, Exp-Audit shows that it is possible to achieve a significantly smaller cost under suitable assumptions. This motivates the study of scenarios where the audit cost is better than that for the black-box case, as an interesting future direction.
@@
The design of Exp-Audit has two novel aspects. The first is to view the past database under partial feedback as a collection of truncated samples from a distribution which allows us to use the past data for learning the distribution parameters using techniques for learning from truncated samples LWZ (23); DGTZ (18). The second is to use MAP oracles as a proxy for true labels under a suitable separation condition (which is weaker than that required for recovering a good clustering Das (99); DS (07); VW (02); KSV (05), see Mah (11) Table 1). To do this, we generalize the result on learning with MAP oracles from spherical Gaussians Mah (11) to general exponential family which may be of independent interest and show that under the novel cost model introduced in this work, it leads to direct improvement in the cost incurred by the algorithm.
@@
In Exp-Audit, besides solving the fairness audit problem, we also recover the parameters $q_{y\mid a}$ and $\theta^{*}_{y,a}$ within $\varepsilon$ accuracy. With these parameters, one can recover a classifier whose error rate is at most $O(\varepsilon)$ more than the error rate of the Bayes optimal classifier. This also allows us to study the fairness-accuracy tradeoff since we can now express the Bayes error and unfairness metrics using distribution parameters with high accuracy CDPF+ (17); DWY+ (20); ZG (22).
@@ 6 Experiments & Results
Figure 1: From left, the first two plots show the comparison of the Baseline (Algorithm 3, Appendix A) and the RS-Audit (Algorithm 1) on the Adult Income dataset. The last two plots show the comparison of RS-Audit and Exp-Audit (Algorithm 2).
@@
In the interest of space, a more exhaustive set of experiments are presented in Appendix D. The code can be found at https://github.com/nirjhar-das/fairness_audit_partial_feedback. Here we only present some key results.
@@
Blackbox Model: We compare the Baseline (Algorithm 3) and RS-Audit algorithms on the Adult Income BK (96) and Law School Admission WRC (98) datasets (details in Appendix D). For a particular dataset, we first train a classifier on a portion of the dataset. Thereafter, we run our algorithms with this classifier. We simulate past database using a portion of the dataset that is classified positively by the trained classifier, and online samples by drawing random data points from the dataset. We repeat the experiments over $5$ seeds and plot the average and the standard deviations. The details of the experimental set up are presented in Appendix D. In our experiments, we vary the stopping threshold $\tau$ over a range and study the behavior of the two algorithms. For every $\tau$, we obtain the mean number of samples, the mean predicted EOD and the mean cost. In Fig. 1 (first and second plots from left), we show the results for a logistic regression classifier trained on all the features on the Adult Dataset. The first plot shows the predicted EOD vs average number of samples collected by the two algorithms. We observe that RS-Audit is able to estimate the true EOD (horizontal solid line) with much less samples than Baseline. The second plot shows that RS-Audit’s cost is significantly smaller than Baseline at various values of $\tau$.
@@
Mixture Model: In the mixture model, we perform synthetic experiments. We choose spherical Gaussians as our exponential family and ensure the separation condition (1) during construction. We generate samples from this mixture model and first train a logistic regression classifier. Thereafter, we simulate past data by collecting some samples on which the classifier’s prediction is positive. Since we have the parameters of the Gaussian and the mixture weights, the online data is easily simulated. We experiment with varying values of $\varepsilon$ (with $\delta=0.01$) and repeat the experiments over $5$ seeds. In Fig. 1 (third and fourth plot from left), we show the comparison between RS-Audit and Exp-Audit (Algorithm 2). The third plot shows mean EOD predicted by the two algorithms and the true EOD (solid green line) at different values of $\varepsilon$. We see that the EOD prediced by RS-Audit is closer to the true EOD but has a high standard deviation. On the other hand, Exp-Audit underestimates the true EOD by about $0.05$ but has significantly lower standard deviation. For the fourth plot, we set $c_{feat}=0$ and $c_{lab}=1$ (other values shown in Appendix D) and we show the mean audit cost of the two algorithms at different values of $\varepsilon$. We observe that the audit cost of Exp-Audit is significantly smaller than the audit cost of RS-Audit. Moreover, the trend emerges that the cost of Exp-Audit is independent of $\varepsilon$ while the cost of RS-Audit increases with $\varepsilon$ (note that we had to cap $\tau$ for RS-Audit to a maximum of $1000$ for timely execution of the code, hence flattening of the curve).

### v2

@@ 4.1 RS-Audit: Rejection Sampling based Audit
$\mathbb{P}[f=1\mid Y=y,A=a]$ $=\frac{\mathbb{P}[f=1,Y=y\mid A=a]}{\mathbb{P}[Y=y\mid A=a]}~.$
@@
1: Input $\varepsilon,\delta\in(0,1)$ and past database $D$, then set $\tau={576\log(8\lvert\mathcal{A}\rvert/\delta)}/{\varepsilon^{2}}$.
@@
4:   $N\leftarrow\texttt{Online}(\tau,y,a)$, $N^{\prime}\leftarrow\texttt{Past}(D,\tau,y,a)$.
5:   Set $\widehat{q}_{y\mid a}=\frac{\tau}{N}$ and $\widehat{p}_{y\mid a}=\frac{\tau}{N^{\prime}}$.
[…]
@@
Although RS-Audit and baseline (Algorithm 3) are superficially similar, these are two conceptually different algorithms. The novel insight of not observing the true labels for individuals with $A\neq a$ when trying to estimate $q_{y\mid a}$ exploits the structure of the audit cost under partial feedback where $A$ is observable without cost, which accrues only when the auditor requests the true labels. Moreover, we crucially identify the right amount of information needed to estimate $\mathbb{P}[f=1\mid Y=y,A=a]$, namely, samples of true labels from the distribution $\mathbb{P}[Y\mid A=a]$. Combining these two observations leads to RS-Audit, which is nearly optimal in cost complexity (see Theorem 3). Moreover, due to these observations, the gap in sample complexity of baseline over RS-Audit is roughly $\widetilde{\Omega}\left(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{\mathbb{P}[f=0,A\neq a]}{q_{y,a}\varepsilon^{2}}\right)$ and the difference in audit cost is roughly $\widetilde{\Omega}\left(\sum_{y\in\{0,1\},a\in{\mathcal{A}}}\frac{c_{feat}\mathbb{P}[f=0,A\neq a]+c_{lab}\mathbb{P}[Y=0,f=0,A\neq a]}{q_{y,a}\varepsilon^{2}}\right)$.
@@ 5 Fairness Audit under Mixture Model
Algorithm 2 Exp-Audit: Audit for Mixture Model
@@
1: Set $\varepsilon^{\prime}=\frac{1}{10}$, $\tau^{\prime}=\frac{4\log(14\lvert{\mathcal{A}}\rvert/\delta)}{\varepsilon^{\prime 2}}$, $\tau=\frac{576\log(14\lvert{\mathcal{A}}\rvert/\delta)}{\varepsilon^{2}}$.
@@
5:   $N^{\prime}\leftarrow\texttt{Online}(\tau^{\prime},y,a)$, $N\leftarrow\texttt{Past}(D,\tau,y,a)$
@@
7:  Create MAP oracle: $M_{a}(x)=\underset{y\in\{0,1\}}{\text{argmax}}\ \widetilde{q}_{y\mid a}\mathcal{E}_{\widehat{\theta}_{y,a}}(x).$
8:  Let $\widetilde{q}_{m,a}=\widetilde{q}_{1\mid a}\wedge\widetilde{q}_{0\mid a},~R_{0}\!=\!R_{1}\!=\!0$. Observe $R=\frac{3430\log(12\lvert{\mathcal{A}}\rvert/\delta)}{\widetilde{q}_{m,a}\varepsilon^{2}}$ features $X$ with A=aA\!=\!a and set $R_{M_{a}(X)}\leftarrow R_{M_{a}(X)}+1$ for each $X$. Set $\widehat{q}_{y\mid a}=\frac{R_{y}}{R}$
@@
[…]
Recall that in the mixture model, given $A=a$, the true label is sampled according to $Y\sim\mathrm{Ber}(q_{1\mid a})$, and given both $A=a$ and $Y=y$, the feature of the individual is sampled according to $X\sim{\mathcal{E}}_{\theta^{*}_{y,a}}$ (see Section 23). Also, we denote $\mathbf{c}\coloneqq(L,\beta,\kappa,\alpha)$, which are some known constants of the problem. Now we describe Exp-Audit (Algorithm 2) solving fairness audit under partial feedback in this model with a significantly lower audit cost than the blackbox case.
@@
Exp-Audit first splits the past database of positively classified individuals into subsets $D_{y,a}$, $a\in{\mathcal{A}}$, $y\in\{0,1\}$, so that $D_{y,a}$ only contains individuals with $A=a$ and $Y=y$. Note that samples in $D_{y,a}$ are iid samples of $X\sim{\mathcal{E}}_{\theta^{*}_{y,a}}$ truncated by $f$, i.e., $D_{y,a}$ contains only those samples from ${\mathcal{E}}_{\theta^{*}_{y,a}}$ for whom $f=1$. On every subset $D_{y,a}$, Exp-Audit applies the subroutine $\texttt{TruncEst}(D_{y,a},f,\frac{\varepsilon}{2\beta},\frac{\delta}{14\lvert{\mathcal{A}}\rvert},\mathbf{c},\Theta)$ (adaptation of Lee et al. (2023) and polynomial time, see Appendix B.2). From $D_{y,a}$ this subroutine learns an estimate $\widehat{\theta}_{y,a}$ of $\theta^{*}_{y,a}$ such that $\lVert\widehat{\theta}_{y,a}-\theta^{*}_{y,a}\rVert\leq\frac{\varepsilon}{2\beta}$ with probability at least $1-\frac{\delta}{14\lvert{\mathcal{A}}\rvert}$ if $\lvert D_{y,a}\rvert\geq\widetilde{\Omega}(\frac{d}{\varepsilon^{2}})$. Note that the condition $\lvert D_{y,a}\rvert\geq\widetilde{\Omega}(\frac{d}{\varepsilon^{2}})$ can be assumed true without loss of generality because of Assumption 1(i) which states $\mathbb{P}[f=1\mid Y=y,A=a]>0$, ensuring that the condition is satisfied in finite time. Crucially, there is no cost for waiting for samples which being iid are unaffected by delay.
After obtaining $\widehat{\theta}_{y,a}$, the algorithm uses the rejection-sampling based approach (line 5) similar to RS-Audit to obtain $\widehat{p}_{y\mid a}\in(1\pm\frac{\varepsilon}{12})p_{y\mid a}$ and $\widetilde{q}_{y\mid a}\in(1\pm\varepsilon^{\prime})q_{y\mid a}$, where $\varepsilon^{\prime}=\frac{1}{10}$ is a constant. Note that the cost incurred in estimating $\widetilde{q}_{y,a}$ is therefore $\widetilde{O}(1)$. This is a crucial difference of Exp-Audit with RS-Audit which estimates $q_{y\mid a}$ upto a multiplicative $1\pm\varepsilon$ error incurring $\widetilde{O}(\frac{1}{\varepsilon^{2}})$ cost.
@@
Then, for every $a\in{\mathcal{A}}$, Exp-Audit constructs a maximum a posteriori (MAP) oracle for the mixture model $M_{a}:{\mathcal{X}}\to\{0,1\}$, $M_{a}(X)=\operatorname*{arg\,max}_{y\in\{0,1\}}\widetilde{q}_{y\mid a}{\mathcal{E}}_{\widehat{\theta}_{y,a}}(X)$. In other words, for $A=a$ and feature $X$, the MAP oracle returns which is the most likely true label $Y$. Next, for every $a\in{\mathcal{A}}$, with $\widetilde{q}_{m,a}=\widetilde{q}_{0\mid a}\wedge\widetilde{q}_{1\mid a}$, the algorithm observes only the features $X$ for $R=\widetilde{O}(\frac{1}{\widetilde{q}_{m,a}\varepsilon^{2}})$ individuals from group $A=a$ and uses the MAP oracle $M_{a}$ to assign a proxy label to each $X$. With $R_{y}$ as the number of individuals with proxy label $y$, Exp-Audit computes $\widehat{q}_{y\mid a}=\frac{R_{y}}{R}$. Under the separation condition (1), we have $\widehat{q}_{y\mid a}\in(1\pm\frac{\varepsilon}{12})q_{y\mid a}$ (see Appendix B.1). Using $\widehat{p}_{y\mid a}$ and $\widehat{q}_{y\mid a}$, Exp-Audit sets $\widehat{\Delta}=\max_{y\in\{0,1\},a,a^{\prime}\in{\mathcal{A}}}\left\lvert\frac{\widehat{p}_{y\mid a}}{\widehat{q}_{y\mid a}}-\frac{\widehat{p}_{y\mid a^{\prime}}}{\widehat{q}_{y\mid a^{\prime}}}\right\rvert$ and returns $\mathsf{UNFAIR}$ if $\widehat{\Delta}>\gamma+\frac{\varepsilon}{2}$, else $\mathsf{FAIR}$.
@@
The cost incurred by Exp-Audit is significantly lower than that of RS-Audit (see Theorems 2 and 4) because the term involved with $c_{lab}$, the cost of requesting a negative label ($Y=0$), is a problem-dependent constant and independent of $\varepsilon$ for Exp-Audit. Note that in most applications, $c_{lab}\gg c_{feat}$. This motivates the study of scenarios where the audit cost is lower than the black-box case, as an interesting future direction.
@@
The design of Exp-Audit has two novel aspects. The first is to view the past database under partial feedback as a collection of truncated samples which allows us to learn the distribution parameters using techniques for learning from truncated samples Lee et al. (2023); Daskalakis et al. (2018) using past data. The second is to construct approximate MAP oracles as a proxy for true labels under a suitable separation condition. For this, we generalize the result of Mahalanabis (2011) from spherical Gaussians to general exponential family (may be of independent interest), leading to improvement in the audit cost.
@@
In Exp-Audit, besides solving the audit problem, we also recover the parameters $q_{y\mid a}$ and $\theta^{*}_{y,a}$ within $\varepsilon$ error which can be used to obtain a classifier with error rate at most $O(\varepsilon)$ more than that of the Bayes optimal classifier. Moreover, fairness-accuracy tradeoff can be studied by expressing the Bayes error and unfairness metrics using learnt distribution parameters with high accuracy Corbett-Davies et al. (2017); Dutta et al. (2020); Zhao and Gordon (2022).
Remark 5 (Separation Condition).
The separation condition (1) assumed to obtain Theorem 4 is significantly milder than what is required to obtain good clustering of a mixture of exponential family. We require that the separation is $\Omega(\sqrt{\log({1/\varepsilon})})$ whereas that required for good clustering is $\Omega(\frac{1}{\varepsilon})$ (see Mahalanabis (2011) Table 1). Moreover, this separation is the minimum required for the Bayes optimal classifier to have low ($O(\varepsilon)$) error rate (see Mahalanabis (2011), Remark 8). We also highlight that the condition (1) can be tested after the parameters $\widehat{\theta}_{y,a}$ and $\widetilde{q}_{y\mid a}$ are learnt (lines 2-7 in Algorithm 2), and if the condition is violated, the auditor can switch to Algorithm 1. In our experiments on the Adult Income dataset, we do not enforce this condition but we observe that Algorithm 2 has significantly lower cost while being able to perform the fairness audit successfully.
@@ 6 Experiments & Results
Figure 1: Top Row: Results for the Baseline (Algorithm 3, Appendix A) and the RS-Audit (Algorithm 1) on the Adult Income dataset. Middle and Bottom Rows: Comparison of RS-Audit and Exp-Audit (Algorithm 2) on the Adult Income (middle row) and synthetic (bottom row) dataset.
@@
In the interest of space, all details and an exhaustive set of experiments are presented in Appendix D, with some key results here. The code is provided in the supplementary.
@@
Blackbox Model: We compare the Baseline (Algorithm 3) and RS-Audit on the Adult Income Becker and Kohavi (1996) and Law School Admission Wightman et al. (1998) datasets (details in Appendix D). For a given dataset, we first train a classifier on a subset. Thereafter, we run our algorithms with this classifier. We simulate past database using a portion of the dataset that is classified positively by the trained classifier, and online samples by drawing random data points from the dataset. We repeat the experiments across $5$ seeds and plot the average and standard deviation. The details of the experimental setup are presented in Appendix D. In our experiments, we set $\gamma=0$ and vary the stopping threshold $\tau$ over a range and study the behavior of the two algorithms. For every $\tau$, we obtain the mean number of samples, the mean predicted equalized odds difference (EOD), and the mean cost. In Fig. 1, we show the results for a logistic regression classifier trained on all the features on the Adult Dataset. The first plot shows the predicted EOD vs the average number of samples collected by the two algorithms. We observe that RS-Audit (black line) is able to estimate the true EOD (horizontal solid line) with much fewer samples than Baseline (red line). The second plot shows that RS-Audit’s cost (black line) is significantly smaller than Baseline (red line) at various values of $\tau$.
@@
Mixture Model: In the mixture model, we perform experiments on synthetic and Adult Income datasets. For the synthetic experiments, we choose spherical Gaussians as our exponential family and ensure the separation condition (1) during construction. For the Adult Income dataset, we first train a neural network to predict the labels, and then use the embedding from the penultimate layer of the network. We do not enforce any structure or separation on the embeddings, instead we instantiate Exp-Audit with the exponential family being multivariate Gaussian. Note that the true data may not actually follow this distribution. Moreover, the past data and online samples are simulated similarly as in the experiments for the blackbox model. We experiment with various values of $\gamma$ and $\varepsilon$ (with $\delta=0.01$) and repeat the experiments over $5$ seeds. In Fig. 1, we show the comparison between RS-Audit and Exp-Audit (Algorithm 2). We observe that on both Adult dataset (second column) and the synthetic dataset (third column), while the estimated EOD of RS-Audit is slightly more closer to the true EOD than Exp-Audit, the cost of Exp-Audit is significantly smaller than that of RS-Audit (here we show for $c_{feat}=0$ and $c_{lab}=1$, while other values are shown in Appendix D). Moreover, the trend emerges that the cost of Exp-Audit is independent of $\varepsilon$ while the cost of RS-Audit increases with $\varepsilon$ (note that we had to cap $\tau$ for RS-Audit to a maximum of $1000$ for timely execution of the code, hence flattening of the curve).
Remark 6 (Empirical Choices).
Our algorithms and their guarantees are agnostic to the choice of classifier. Further, the blackbox setting is agnostic to the data distribution. Therefore, we choose logistic regression as our classifier which is also preferred in fairness applications due to its explainability over complicated models, and we experiment with popular fairness datasets to validate our theoretical results empirically. Note that in the mixture model experiment with Adult dataset, we do not enforce any of the structural or separation assumptions required to prove Theorem 4. Yet, we observe that Exp-Audit performs better than RS-Audit in terms of the audit cost while giving good EOD estimates, further indicating its robustness.